# SHIELD: Safety on Humanoids via CBFs In Expectation on Learned Dynamics

### v1

@@ I Introduction
Contributions. This paper introduces SHIELD, a novel paradigm for guaranteeing safety in robotic systems that bridges the gap between data-driven and model-based safety methods. SHIELD is specifically designed for systems with complex, robust, but ultimately stochastic low-level controllers, such as RL policies used by humanoid robots for locomotion. Unlike traditional safety filters, SHIELD functions as a safety layer that sits ”above” the nominal learning-based controller in the autonomy stack (cf. Fig. 1), modulating the reference signal rather than directly filtering control outputs. SHIELD is constructed through a three-step process:
@@
where $\mathbf{F}(\mathbf{x}_{k})+\mathbf{G}(\mathbf{x}_{k})\mathbf{u}_{k}$ represents a simplified model of the system and $\mathbf{d}$ is the difference between the full-order model and this reduced order model, also called the dynamics residual. To capture the complexities of the full-order dynamics $\boldsymbol{\Phi}$ and the controller $\boldsymbol{\pi}$, we consider $\mathbf{d}_{k}$ to be a random disturbance sampled from a distribution $\mathcal{D}(\mathbf{s}_{k:0},\mathbf{a}_{k:0})$ that is dependent on the history of full states and from time 00 through $k$, denoted as $\mathbf{s}_{k:0}$ and $\mathbf{a}_{k:0}$ respectively.
@@ III Disturbance Learning
Conditional Variational Inference. To account for the dynamics residual, we seek to train a generative model to approximate the dynamics residual distribution. To do this, we first collect a dataset of state, command, and disturbance tuples $\mathfrak{D}=\{(\mathbf{x}_{i},\mathbf{u}_{i},\mathbf{d}_{i})\}_{i=1}^{N}$. We then train a Conditional Variation Autoencoder (CVAE) [30] on this dataset, which yields a generative disturbance model $p_{\theta}(\mathbf{d}_{k}|\mathbf{x}_{k:k-N},\mathbf{u}_{k:k-N})$. In contrast to previous work [23], the model is conditioned on a context window of length $N\in\mathbb{N}$, to allow the model to better capture temporal effects such as higher state derivatives or time delays. We find that providing this context greatly boosts modeling accuracy for a complex system like a humanoid robot (Sec. V). Note that the input $\mathbf{u}_{i}$ here is the unfiltered command, meaning we do not need to solve the algebraic loop of the filtered input (III) being a function of itself or its own history.
@@
Figure 3: Higher $\alpha=L(P,K=10,h(x_{0})=10,\delta=0.01,\sigma)$ values correspond to more conservative behavior, this increased conservatism a consequence of a lower $K$-step exit probability or a higher variance.
@@
In practice we approximate $L:(0,1)\times\mathbb{N}\times\mathbb{R}_{>0}\times\mathbb{R}_{>0}\times\mathbb{R}_{>0}\to(0,1)$ numerically due to the complexity of the analytic solution. Evaluations of the function for different $\alpha$ values for a range of $P$ and $\sigma$ can be found in Fig. 3.
@@ IV Dynamic obstacle avoidance on stochastic reduced-order models
Figure 6: SHIELD enforces safety in collision avoidance with adaptive conservatism. The A* planner path is not necessarily safe even it does not cross the obstacle, thus naively following the path would result in collisions or scrapes. Nominal CBF, due to not accounting for the inaccurate reduced order model, would also result in collisions or be extremely conservative.
@@ V Experiments
$P_{\textup{failure}}=\frac{\sum_{i=1}^{N_{\textup{trials}}}\sum_{j=1}^{N_{\textup{steps}}}\mathbbm{1}_{\{h(\mathbf{x}_{i,j})<0}\}}{N_{\textup{trials}}N_{\textup{steps}}}$ (25)
@@
Obstacle Avoidance. First, we conduct controlled experiments with fixed obstacles. We define success as the robot walking past obstacles without making contact. We model the detected obstacles as cylinders of radius 0.3m and the robot to have a safety margin of 0.38m from the center of mass. To navigate, we first use A* [39] to first plan a path through free space, we then generate nominal velocities by directing the robot from its current position to the next node on the path and filter the commanded velocities with SHIELD. We present both single-obstacle and multi-obstacle cases. In single-obstacle experiments, naively following the A* path alone does not completely avoid obstacles due to state tracking errors. The nominal DTCBF filter, being unaware of the dynamics residual, either collides into the obstacle or exhibits extreme conservative behavior with $\alpha=0.99$. However, SHIELD enables the robot to completely bypass the obstacle. We observe similar behavior in multi-obstacle scenarios, where SHIELD is able to adjust conservativeness online to only enforce maximum safety conditions when needed, resulting in more dynamic behavior. The results of these experiments can be seen in Fig. 6.

### v2

@@ I Introduction
Contributions. This paper introduces SHIELD, a novel paradigm for guaranteeing safety in robotic systems that bridges the gap between data-driven and model-based safety methods. SHIELD is specifically designed for systems with complex, robust, but ultimately stochastic low-level controllers, such as RL policies used by humanoid robots for locomotion. Unlike traditional safety filters, SHIELD functions as a safety layer that sits “above” the nominal learning-based controller in the autonomy stack (cf. Fig. 1), modulating the reference signal rather than directly filtering control outputs. SHIELD is constructed through a three-step process:
@@
where $\mathbf{F}(\mathbf{x}_{k})+\mathbf{G}(\mathbf{x}_{k})\mathbf{u}_{k}$ represents a simplified model of the system and $\mathbf{d}$ is the difference between the full-order model and this reduced order model, also called the dynamics residual. To capture the complexities of the full-order dynamics $\boldsymbol{\Phi}$ and the controller $\boldsymbol{\pi}$, we consider $\mathbf{d}_{k}$ to be a random disturbance sampled from a distribution $\mathcal{D}(\mathbf{s}_{k:0},\mathbf{a}_{k:0})$ that is dependent on the history of full states and from time 0 through $k$, denoted as $\mathbf{s}_{k:0}$ and $\mathbf{a}_{k:0}$ respectively.
@@ III Disturbance Learning
Conditional Variational Inference. To account for the dynamics residual, we seek to train a generative model to approximate the dynamics residual distribution. To do this, we first collect a dataset of state, command, and disturbance tuples $\mathfrak{D}=\{(\mathbf{x}_{i},\mathbf{u}_{i},\mathbf{d}_{i})\}_{i=1}^{N}$. We then train a Conditional Variation Autoencoder (CVAE) [30] on this dataset with latent size 3 which yields a generative disturbance model $p_{\theta}(\mathbf{d}_{k}|\mathbf{x}_{k:k-N},\mathbf{u}_{k:k-N})$. In contrast to previous work [23], the model is conditioned on a context window of length $N\in\mathbb{N}$, to allow the model to better capture temporal effects such as higher state derivatives or time delays. We find that providing this context greatly boosts modeling accuracy for a complex system like a humanoid robot (Sec. V). Note that the input $\mathbf{u}_{i}$ here is the unfiltered command, meaning we do not need to solve the algebraic loop of the filtered input (III).
@@
Figure 3: Higher $\alpha=L(P,K=10,h(x_{0})=10,\delta=0.01,\sigma)$ values correspond to more conservative behavior. This increased conservatism is a consequence of a lower $K$-step exit probability or a higher variance.
@@
In practice we approximate $L:(0,1)\times\mathbb{N}\times\mathbb{R}_{>0}\times\mathbb{R}_{>0}\times\mathbb{R}_{>0}\to(0,1)$ numerically by performing root finding with Brent’s Method [35] due to the complexity of the analytic solution. Evaluations of the function for different $\alpha$ values for a range of $P$ and $\sigma$ can be found in Fig. 3.
@@ IV Dynamic obstacle avoidance on stochastic reduced-order models
Figure 6: SHIELD enforces safety in collision avoidance with adaptive conservatism. The A* planner path is not necessarily safe even though it does not cross the obstacle, thus naively following the path would result in collisions or scrapes. Nominal CBF, due to not accounting for the inaccurate reduced order model, would also result in collisions or be extremely conservative.
@@ V Experiments
$P_{\textup{failure}}=\frac{\sum_{i=1}^{N_{\textup{trials}}}\sum_{j=1}^{N_{\textup{steps}}}\mathds{1}_{\{h(\mathbf{x}_{i,j})<0\}}}{N_{\textup{trials}}N_{\textup{steps}}}$ (25)
@@
Obstacle Avoidance. First, we conduct controlled experiments with fixed obstacles. We define success as the robot walking past obstacles without making contact. We model the detected obstacles as cylinders of radius 0.3m and the robot to have a safety margin of 0.38m from the center of mass. To navigate, we first use A* [40] to first plan a path through free space, we then generate nominal velocities by directing the robot from its current position to the next node on the path and filter the commanded velocities with SHIELD. We present both single-obstacle and multi-obstacle cases. In single-obstacle experiments, naively following the A* path alone does not completely avoid obstacles due to state tracking errors. The nominal DTCBF filter, being unaware of the dynamics residual, either performs too aggressively colliding with the obstacle, or performs to conservatively taking almost twice the time to arrive at the goal. However, SHIELD enables the robot to completely bypass the obstacle. We observe similar behavior in multi-obstacle scenarios, where SHIELD is able to adjust conservativeness online to only enforce maximum safety conditions when needed, resulting in more dynamic behavior. The results of these experiments can be seen in Fig. 6.